# Resurgent Wilson loops in refined topological string

### v1

@@ 4.4 Example: local $\mathbb{P}^{1}\times\mathbb{P}^{1}$
Similar to Section 4.3, we consider the example of refined topological string theory on local $\mathbb{P}^{1}\times\mathbb{P}^{1}$, i.e. the total space of the canonical bundle of $\mathbb{P}^{1}\times\mathbb{P}^{1}$, with the constraint that the two $\mathbb{P}^{1}$s have the same volume, also known as the massless limit. This theory has also been discussed in detail in the literature. It has a one dimensional moduli space with three singular points of large radius, conifold, and orbifold types Haghighat:2008gw , and we take the convention that it is parametrised by the global complex coordinate $z$, such that the three singular points are located at $z=0$, $z=1/16$, and $z=\infty$ respectively [^13] Note that in the case of massless local $\mathbb{F}_{0}$, the orbifold point at $z=\infty$ has a conifold singularity superimposed upon it so that the free energies satisfy the gap conditions at both the conifold point and the orbifold point..
@@
Near the large radious point, the flat coordinate and its conjugate are (see e.g. Marino:2015ixa )
@@
As in Section 4.3, we first inspect the non-perturbative corrections for Wilson loop BPS sectors, which can be calculated effectively using the algorithm in Wang:2023zcb . For simplicity, we focus on the range between the large radious point $z=0$ and the conifold point $z=1/16$.
@@
Figure 4.14: Borel singularities of refined Wilson loop BPS sectors $\mathcal{F}[1](\mathsf{b};g_{s})$ for local $\mathbb{P}^{1}\times\mathbb{P}^{1}$ with $\mathsf{b}=2$ up to $g=50$ in the large radius frame, respectively (a) near the large radius point $z=0$ and (b) near the conifold point $z=1/16$. The red dots are the approximate singularities from numerical calculations, and the black dots are the indicated branch points, located at (a) $\mathsf{b}^{-1}\mathcal{A}_{(-2,0,0)_{\text{LR}}},\mathsf{b}^{-1}\mathcal{A}_{(-2,1,0)_{\text{LR}}},\mathsf{b}^{-1}\mathcal{A}_{(-2,2,-1)_{\text{LR}}},\mathsf{b}^{-1}\mathcal{A}_{(-2,3,-2)_{\text{LR}}}$ and (b) $\mathsf{b}^{-1}\mathcal{A}_{(-2,0,0)_{\text{LR}}}$ respectively.
@@
Figure 4.15: Borel singularities of refined Wilson loop BPS sectors $\mathcal{F}[2](\mathsf{b};g_{s})$ for local $\mathbb{P}^{1}\times\mathbb{P}^{1}$ with $\mathsf{b}=2$ up to $g=50$ in the large radius frame, respectively (a) near the large radius point $z=0$ and (b) near the conifold point $z=1/16$. The red dots are the approximate singularities from numerical calculations, and the black dots are the indicated branch points, located at (a) $\mathsf{b}^{-1}\mathcal{A}_{(-2,0,0)_{\text{LR}}},\mathsf{b}^{-1}\mathcal{A}_{(-2,1,0)_{\text{LR}}},\mathsf{b}^{-1}\mathcal{A}_{(-2,2,-1)_{\text{LR}}},\mathsf{b}^{-1}\mathcal{A}_{(-2,3,-2)_{\text{LR}}}$ and (b) $\mathsf{b}^{-1}\mathcal{A}_{(-2,0,0)_{\text{LR}}}$ respectively.
@@
Figure 4.16: Borel singularities of refined free energies $\mathcal{F}[0](\mathsf{b};g_{s})$ for local $\mathbb{P}^{1}\times\mathbb{P}^{1}$ with $\mathsf{b}=2$ up to $g=50$ in the large radius frame, respectively (a) near the large radius point $z=0$ and (b) near the conifold point $z=1/16$. The red dots are the approximate singularities from numerical calculations, and the black dots are the indicated branch points, located at (a) $\mathsf{b}^{-1}\mathcal{A}_{(0,1,n)_{\text{LR}}}$ ($n=0,1,2$) and (b) $\mathsf{b}^{-1}\mathcal{A}_{(-2,0,0)_{\text{LR}}}$ respectively.
Let us study the location of Borel singularities first. We evaluate the perturbative BPS sectors $\mathcal{F}[1]$ and $\mathcal{F}[2]$ in the holomorphic limit of the large radius frame, where $t_{\text{LR}}$ is the flat coordinate, near respectively the large radius point $z=0$ and the conifold point $z=1/16$. The Borel singularities of $\mathcal{F}[1]$ and $\mathcal{F}[2]$ are plotted in Figs. 4.14 and Figs. 4.15. These two plots are similar. Near the large radius point, the visible Borel singularities are located at $\mathsf{b}^{-1}\mathcal{A}_{\gamma_{\text{LR}}}$ (we take $\mathsf{b}>1$ so that $\mathsf{b}^{-1}\mathcal{A}$ is smaller than $\mathsf{b}\mathcal{A}$) with the charges
@@
Near the conifold point, the visible Borel singularities are located at $\mathsf{b}^{-1}\mathcal{A}_{\pm(-2,0,0)_{\text{LR}}}$. For comparison, we also give the same plots for the free energies [^14] The constant map contributions to free energies are removed. $\mathcal{F}[0]$ in Figs. 4.16. The visible Borel singularities are located at $\mathsf{b}^{-1}\mathcal{A}_{\gamma_{\text{LR}}}$ with
@@
Figure 4.17: Borel singularities of refined Wilson loop BPS sectors $\mathcal{F}[1](\mathsf{b};g_{s})$ for local $\mathbb{P}^{1}\times\mathbb{P}^{1}$ with $\mathsf{b}=2$ up to $g=50$ in the conifold frame, respectively (a) very close to the large radius point and (b) away from it toward the conifold point. The red dots are the approximate singularities from numerical calculations, and the black dots are the indicated branch points, located at (a) $\mathsf{b}^{-1}\mathcal{A}_{(0,1,n)_{\text{LR}}}$ ($n=0,1,2$) and (b) $\mathsf{b}^{-1}\mathcal{A}_{(-2,0,0)_{\text{c}}},\mathsf{b}^{-1}\mathcal{A}_{(-2,1,0)_{\text{c}}},\mathsf{b}^{-1}\mathcal{A}_{(-2,0,1)_{\text{c}}}$ respectively.
@@
Figure 4.18: Borel singularities of refined Wilson loop BPS sectors $\mathcal{F}[2](\mathsf{b};g_{s})$ for local $\mathbb{P}^{1}\times\mathbb{P}^{1}$ with $\mathsf{b}=2$ up to $g=50$ in the conifold frame, respectively (a) very close to the large radius point and (b) away from it toward the conifold point. The red dots are the approximate singularities from numerical calculations, and the black dots are the indicated branch points, located at (a) $\mathsf{b}^{-1}\mathcal{A}_{(0,1,n)_{\text{LR}}}$ ($n=0,1,2$) and (b) $\mathsf{b}^{-1}\mathcal{A}_{(-2,0,0)_{\text{c}}},\mathsf{b}^{-1}\mathcal{A}_{(-2,1,0)_{\text{c}}},\mathsf{b}^{-1}\mathcal{A}_{(-2,0,1)_{\text{c}}}$ respectively.
@@
Next, we study the non-perturbative series. We focus on the 1-instanton sector as Section in 4.4. The large order asymptotics of the perturbative coefficients are the same as (4.64) and (4.66). We consider two cases, the BPS sectors near the conifold point in the large radius frame, as well as in the conifold frame. The dominant Borel singularities are respectively $\gamma_{\text{LR}}=\pm(-2,0,0)$ and $\gamma_{\text{c}}=\pm(-2,0,0)$, as shown in the plots of Figs. 4.14 (b), 4.15 (b) and Figs. 4.17 (b), 4.18 (b). we compare these numerical results of $\mu_{0},\mu_{1},\mu_{2},\ldots$ extracted from perturbative data using the large order formulas with the theoretical prediction from Sections 4.1 and 4.2.1 in Figs. 4.19, 4.20, 4.21, 4.22 for generic $\mathsf{b}$, and in Figs. 4.23, 4.24, 4.25, 4.26 for the unrefined limit $\mathsf{b}=1$.

### v2

@@ 5.2 Example: local $\mathbb{P}^{1}\times\mathbb{P}^{1}$
Similar to Section 5.1, we consider the example of refined topological string theory on local $\mathbb{P}^{1}\times\mathbb{P}^{1}$, i.e. the total space of the canonical bundle of $\mathbb{P}^{1}\times\mathbb{P}^{1}$, with the constraint that the two $\mathbb{P}^{1}$s have the same volume, also known as the massless limit. This theory has also been discussed in detail in the literature. It has a one dimensional moduli space with three singular points of large radius, conifold, and orbifold types Haghighat:2008gw , and we take the convention that it is parametrized by the global complex coordinate $z$, such that the three singular points are located at $z=0$, $z=1/16$, and $z=\infty$ respectively [^13] Note that in the case of massless local $\mathbb{F}_{0}$, the orbifold point at $z=\infty$ has a conifold singularity superimposed upon it so that the free energies satisfy the gap conditions at both the conifold point and the orbifold point..
@@
Near the large radius point, the flat coordinate and its conjugate are (see e.g. Marino:2015ixa )
@@
As in Section 5.1, we first inspect the non-perturbative corrections for Wilson loop BPS sectors, which can be calculated effectively using the algorithm in Wang:2023zcb . For simplicity, we focus on the range between the large radius point $z=0$ and the conifold point $z=1/16$.
@@
Figure 5.14: Borel singularities of refined Wilson loop BPS sectors $\mathcal{F}[1](\mathsf{b};g_{s})$ for local $\mathbb{P}^{1}\times\mathbb{P}^{1}$ with $\mathsf{b}=2$ up to $g=50$ in the large radius frame, respectively (a) near the large radius point $z=0$ and (b) near the conifold point $z=1/16$. The red dots are approximate singularities from numerical calculations, which would accumulate to branch cuts. The branch points (black dots) on the imaginary axis are $\mathsf{b}^{-1}\mathcal{A}_{\pm(-2,0,0)_{\text{LR}}}$, and those away from the imaginary axis are $\mathsf{b}^{-1}\mathcal{A}_{\pm(-2,1,0)_{\text{LR}}},\mathsf{b}^{-1}\mathcal{A}_{\pm(-2,2,-1)_{\text{LR}}},\mathsf{b}^{-1}\mathcal{A}_{\pm(-2,3,-2)_{\text{LR}}}$.
@@
Figure 5.15: Borel singularities of refined Wilson loop BPS sectors $\mathcal{F}[2](\mathsf{b};g_{s})$ for local $\mathbb{P}^{1}\times\mathbb{P}^{1}$ with $\mathsf{b}=2$ up to $g=50$ in the large radius frame, respectively (a) near the large radius point $z=0$ and (b) near the conifold point $z=1/16$. The red dots are approximate singularities from numerical calculations, which would accumulate to branch cuts. The branch points (black dots) are the same as in Fig. 5.14.
@@
Figure 5.16: Borel singularities of refined free energies $\mathcal{F}[0](\mathsf{b};g_{s})$ for local $\mathbb{P}^{1}\times\mathbb{P}^{1}$ with $\mathsf{b}=2$ up to $g=50$ in the large radius frame, respectively (a) near the large radius point $z=0$ and (b) near the conifold point $z=1/16$. The red dots are approximate singularities from numerical calculations. The branch points (black dots) are (a) $\mathsf{b}^{-1}\mathcal{A}_{\pm(0,1,n)_{\text{LR}}}$ ($n=-1,-2,0,1,2$) and (b) $\mathsf{b}^{-1}\mathcal{A}_{\pm(-2,0,0)_{\text{LR}}}$ respectively.
Let us study the location of Borel singularities first. We evaluate the perturbative BPS sectors $\mathcal{F}[1]$ and $\mathcal{F}[2]$ in the holomorphic limit of the large radius frame, where $t_{\text{LR}}$ is the flat coordinate, near respectively the large radius point $z=0$ and the conifold point $z=1/16$. The Borel singularities of $\mathcal{F}[1]$ and $\mathcal{F}[2]$ are plotted in Fig. 5.14 and Fig. 5.15. These two plots are similar. Near the large radius point, the visible Borel singularities are located at $\mathsf{b}^{-1}\mathcal{A}_{\gamma_{\text{LR}}}$ (we take $\mathsf{b}>1$ so that $\mathsf{b}^{-1}\mathcal{A}$ is smaller than $\mathsf{b}\mathcal{A}$) with the charges
@@
Near the conifold point, the visible Borel singularities are located at $\mathsf{b}^{-1}\mathcal{A}_{\pm(-2,0,0)_{\text{LR}}}$. For comparison, we also give the same plots for the free energies [^14] The constant map contributions to free energies are removed. $\mathcal{F}[0]$ in Fig. 5.16. The visible Borel singularities are located at $\mathsf{b}^{-1}\mathcal{A}_{\gamma_{\text{LR}}}$ with
@@
Figure 5.17: Borel singularities of refined Wilson loop BPS sectors $\mathcal{F}[1](\mathsf{b};g_{s})$ for local $\mathbb{P}^{1}\times\mathbb{P}^{1}$ with $\mathsf{b}=2$ up to $g=50$ in the conifold frame, respectively (a) very close to the large radius point and (b) away from it toward the conifold point. The red dots are approximate singularities from numerical calculations. The branch points (black dots) are (a) $\mathsf{b}^{-1}\mathcal{A}_{\pm(0,1,n)_{\text{LR}}}$ ($n=-2,-1,0,1,2$) and (b) $\mathsf{b}^{-1}\mathcal{A}_{\pm(-2,0,0)_{\text{c}}}$ (horizontal), $\mathsf{b}^{-1}\mathcal{A}_{\pm(-2,\pm 1,0)_{\text{c}}}$ (slight away), $\mathsf{b}^{-1}\mathcal{A}_{\pm(-2,0,\pm 1)_{\text{c}}}$ (further away) respectively.
@@
Figure 5.18: Borel singularities of refined Wilson loop BPS sectors $\mathcal{F}[2](\mathsf{b};g_{s})$ for local $\mathbb{P}^{1}\times\mathbb{P}^{1}$ with $\mathsf{b}=2$ up to $g=50$ in the conifold frame, respectively (a) very close to the large radius point and (b) away from it toward the conifold point. The red dots are approximate singularities from numerical calculations. The branch points (black dots) are the same as in Fig. 5.17.
@@
Next, we study the non-perturbative series. We focus on the 1-instanton sector as Section in 5.2. The large order asymptotics of the perturbative coefficients are the same as (5.13) and (5.15). We consider two cases, the BPS sectors near the conifold point in the large radius frame, as well as in the conifold frame. The dominant Borel singularities are respectively $\gamma_{\text{LR}}=\pm(-2,0,0)$ and $\gamma_{\text{c}}=\pm(-2,0,0)$, as shown in the plots of Figs. 5.14 (b), 5.15 (b) and Figs. 5.17 (b), 5.18 (b). we compare these numerical results of $\mu_{0},\mu_{1},\mu_{2},\ldots$ extracted from perturbative data using the large order formulas with the theoretical prediction from Sections 4.1 and 4.2.1 in Figs. 5.19, 5.20, 5.21, 5.22 for generic $\mathsf{b}$, and in Figs. 5.23, 5.24, 5.25, 5.26 for the unrefined limit $\mathsf{b}=1$.